# Diffusion-based Extreme Image Compression with Compressed Feature Initialization

### v1

@@ 4.2 Experimental Results
Table 3: Ablation studies of Relay Residual Diffusion (RRD) framework and Fixed-Step Fine-Tuning (FSFT) strategy. BD-rate is calculated on CLIC2020 dataset. DS denotes denoising steps.
Methods Starting Point DS w/ RRD w/ FSFT BD-rate (%)
Distortion Perception Average
Baseline $\mathcal{N}(0,\bm{I})$ 2/5 $\times$ $\times$ 209.0 247.0 228.0
50 $\times$ $\times$ 100.9 50.7 75.8
v_1 $\bm{z}_{N}$ 2/5 $\times$ 27.2 46.4 36.8
RDEIC (Ours) $\bm{z}_{N}$ 2/5 0 0 0
4.3 Ablations
@@
To provide a more comprehensive analysis of the proposed method, we conduct ablation studies, with the results presented in Table 3. For the baseline, we employ the same diffusion framework as DiffEIC (Li et al., 2024b), where the denoising process starts from pure noise. Due to the inherent characteristics of diffusion models, the baseline requires a large number of denoising steps to achieve optimal performance and exhibits poor distortion performance.
@@
Effectiveness of relay residual diffusion. We first investigate the effectiveness of our proposed relay residual diffusion framework. As shown in Table 3, incorporating the relay residual diffusion framework significantly decreases the denoising steps required for reconstruction meanwhile achieves better reconstruction quality (v_1 vs. Baseline). Specifically, v_1 uses only 2 or 5 denoising steps, while the baseline requires up to 50 steps to achieve similar perceptual performance, indicating that our relay residual diffusion framework can efficiently remove both added noise and residuals. Moreover, v_1 significantly outperforms the baseline in terms of distortion metrics, demonstrating that its reconstruction results have higher fidelity. The reason behind this is that starting from the compressed latent feature, instead of pure noise, avoids the error accumulation in the initial stage of the denoising process and provides a solid foundation for subsequent detail generation.
@@
Analysis for fixed-step fine-tuning. We further demonstrate the effectiveness of the FSFT strategy. Before doing this, we need to analyze the impact of denoising steps in order to select an appropriate value of $L$ for this strategy. As shown in Fig. 6, for $\lambda_{r}\in\{2,1\}$, the number of denoising steps has minimal effect on compression performance, so that we set $L$ to 2 in this case. For $\lambda_{r}\in\{0.5,0.25,0.1\}$, increasing the denoising steps achives better perceptual results (lower LPIPS and DISTS values), but leads to degraded fidelity (lower PSNR and MS-SSIM values). To achieve a balance between fidelity and perceptual quality, we set $L$ to 5 here. In Table 3, using FSFT can yield bitrate savings of 27.2% and 46.4% in terms of distortion and perception metrics, respectively (v_1 vs. RDEIC).
Smoothness-sharpness trade-off. To fully leverage the generative potential of pre-trained stable diffusion, we introduce a controllable detail generation method that allows users to explore and customize outputs according to their personal preferences. The visualization result is shown in Fig. 7. We control the balance between smoothness and sharpness by adjusting the parameter $\lambda_{s}$, which regulates the amount of high-frequency details introduced into the reconstructed image. Specifically, as the value of $\lambda_{s}$ increases, the image transitions from a smooth appearance to a progressively sharper and more detailed reconstruction. Additional results are provided in Fig. 13, Fig. 14, and Fig. 15 in Appendix C.
Figure 6: Quantitative comparisons of different number of denoising steps.
@@
Figure 7: Balancing smoothness versus sharpness. The second row shows the absolute difference between the reconstructed images and the baseline ($\lambda_{s}=0$).
5 Conclusion
@@
Appendix A Mathematical Details
Derivation of Eq. (4). First, according to Eq. (2), $\bm{z}_{n-1}$ can be sampled as:
$\bm{z}_{n-1}$ $=\sqrt{\bar{\alpha}_{n-1}}(\bm{z}_{0}+\eta_{n-1}\bm{e})+\sqrt{1-\bar{\alpha}_{n-1}}\epsilon_{n-1}$ (12)
$=\sqrt{\bar{\alpha}_{n-1}}\bm{z}_{0}+\sqrt{\bar{\alpha}_{n-1}}\eta_{n-1}\bm{e}+\underbrace{\sqrt{1-\bar{\alpha}_{n-1}}\epsilon_{n-1}}_{\sim\mathcal{N}(0,(1-\bar{\alpha}_{n-1})\bm{I})},$ (13)
where $\epsilon_{n-1}\sim\mathcal{N}(0,\bm{I})$. Second, for $\bm{z}_{n}$ defined in Eq. (2) and $\bm{z}_{n-1}$ defined in Eq. (3), we have:
$\bm{z}_{n-1}$ $=k_{n}\bm{z}_{0}+m_{n}\bm{z}_{n}+\sigma_{n}\epsilon$ (14)
$=k_{n}\bm{z}_{0}+m_{n}(\sqrt{\bar{\alpha}_{n}}(\bm{z}_{0}+\eta_{n}\bm{e})+\sqrt{1-\bar{\alpha}_{n}}\epsilon_{n})+\sigma_{n}\epsilon$ (15)
$=(k_{n}+m_{n}\sqrt{\bar{\alpha}_{n}})\bm{z}_{0}+m_{n}\sqrt{\bar{\alpha}_{n}}\eta_{n}\bm{e}+\underbrace{m_{n}\sqrt{1-\bar{\alpha}_{n}}\epsilon_{n}+\sigma_{n}\epsilon}_{\sim\mathcal{N}(0,(m_{n}^{2}(1-\bar{\alpha}_{n})+\sigma_{n}^{2})\bm{I})},$ (16)
where $\epsilon_{n}\sim\mathcal{N}(0,\bm{I})$ and $\epsilon\sim\mathcal{N}(0,\bm{I})$. By combining Eq. (13) and Eq. (16), we obtain the following equations:
Note that, referring to DDIM (Song et al., 2021), we set $\sigma_{n}=0$ for simplicity. By solving Eq. (17), we have:
$k_{n}=\sqrt{\bar{\alpha}_{n-1}}-\sqrt{\frac{1-\bar{\alpha}_{n-1}}{1-\bar{\alpha}_{n}}}\sqrt{\bar{\alpha}_{n}},\ m_{n}=\sqrt{\frac{1-\bar{\alpha}_{n-1}}{1-\bar{\alpha}_{n}}},\ \frac{\eta_{n}}{\eta_{n-1}}=\frac{\sqrt{1-\bar{\alpha}_{n}}/\sqrt{\bar{\alpha}_{n}}}{\sqrt{1-\bar{\alpha}_{n-1}}/\sqrt{\bar{\alpha}_{n-1}}}.$ (18)
Therefore, $\eta_{n}$ can be defined as:
$\eta_{n}=\lambda\frac{\sqrt{1-\bar{\alpha}_{n}}}{\sqrt{\bar{\alpha}_{n}}},$ (19)
where we set $\lambda=\frac{\sqrt{\bar{\alpha}_{N}}}{\sqrt{1-\bar{\alpha}_{N}}}$ to ensure $\eta_{N}=1$.
Derivation of Eq. (8). Substituting Eq. (6) into Eq. (7), we have:
$\|\bm{z}_{0}-\hat{\bm{z}}_{0}\|^{2}_{2}$ $=\|(\frac{\bm{z}_{n}}{\sqrt{\bar{\alpha}_{n}}}-\frac{\sqrt{1-\bar{\alpha}_{n}}}{\sqrt{\bar{\alpha}_{n}}}\tilde{\epsilon}_{n})-(\frac{\bm{z}_{n}}{\sqrt{\bar{\alpha}_{n}}}-\frac{\sqrt{1-\bar{\alpha}_{n}}}{\sqrt{\bar{\alpha}_{n}}}\epsilon_{\theta}(\bm{z}_{n},\bm{c},n))\|^{2}_{2}$ (20)
$=\|\frac{\sqrt{1-\bar{\alpha}_{n}}}{\sqrt{\bar{\alpha}_{n}}}\tilde{\epsilon}_{n}-\frac{\sqrt{1-\bar{\alpha}_{n}}}{\sqrt{\bar{\alpha}_{n}}}\epsilon_{\theta}(\bm{z}_{n},\bm{c},n))\|^{2}_{2}$ (21)
$=\frac{1-\bar{\alpha}_{n}}{\bar{\alpha}_{n}}\|\tilde{\epsilon}_{n}-\epsilon_{\theta}(\bm{z}_{n},\bm{c},n)\|^{2}_{2}$ (22)
Appendix B Additional Experimental Details
Evaluation of third-party models. For VVC (Bross et al., 2021), we used the reference software VTM-23.0222https://vcgit.hhi.fraunhofer.de/jvet/VVCSoftware_VTM/-/tree/VTM-23.0 with intra configuration. The quality factor was selected from the set $\{41,43,45,47,49,52\}$. To compare HiFiC (Mentzer et al., 2020) at extremely low bitrates, we utilized its PyTorch implementation [^3] https://github.com/Justin-Tan/high-fidelity-generative-compression and retrained the model to achieve higher compression ratios, enabling a more direct comparison with our proposed method. For PerCo (Careil et al., 2024), since the official source codes and models are not available, we used a reproduced version [^4] https://github.com/Nikolai10/PerCo/tree/master as a substitute, which employs stable diffusion as the latent diffusion model. For MS-ILLM (Muckley et al., 2023), VQIR (Wei et al., 2024), Text+Sketch (Lei et al., 2023) and DiffEIC (Li et al., 2024b), we used the publicly released checkpoints from their GitHub repositories, and used them for evaluation with the provided code.
Appendix C Additional Experimental Results
Quantitative comparisons on the Tecnick and Kodak datasets. We present the performance of the proposed and compared methods on the Tecnick and Kodak datasets in Fig. 9 and Fig. 10, respectively. The proposed RDEIC achieves state-of-the-art perceptual performance and significantly outperforms other diffusion-based methods in terms of distortion metrics. Since the Kodak dataset is too small to reliably calculate FID and KID scores, we do not report these results for this dataset.
More qualitative results. Fig. 11 and Fig. 12 provide more qualitative comparisons between the proposed RDEIC and other methods. These comparisons reveal the superiority of RDEIC in terms of both fidelity and realism.
Smoothness-sharpness trade-off. As shown in Fig. 13, Fig. 14, and Fig. 15, we control the balance between smoothness and sharpness by adjusting the parameter $\lambda_{s}$, which regulates the amount of high-frequency details introduced into the reconstructed image.
Appendix D Limitations
Ground Truth 0.0271 bpp
Ground Truth 0.0260 bpp
Figure 8: Faces generated at extremely low bitrates.
Figure 9: Quantitative comparisons with state-of-the-art methods on the Tecnick dataset.
Figure 10: Quantitative comparisons with state-of-the-art methods on the Kodak dataset.
Original VVC VQIR MS-ILLM
[…]
Text+Sketch-25 PerCo-20 DiffEIC-50 RDEIC-5
0.0237 / 0.6178 / 0.4237 0.0320 / 0.2130 / 0.7858 0.0278 / 0.1622 / 0.8444 0.0199 / 0.1336 / 0.8699
Original VVC VQIR MS-ILLM
[…]
Text+Sketch-25 PerCo-20 DiffEIC-50 RDEIC-5
0.0163 / 0.2535 / 0.3984 0.0319 / 0.1219 / 0.8874 0.0394 / 0.0664 / 0.9487 0.0391 / 0.0525 / 0.9607
Figure 11: More qualitative comparisons between RDEIC and other methods.
Original VVC VQIR MS-ILLM
[…]
Text+Sketch-25 PerCo-5 DiffEIC-50 RDEIC-2
0.0265 / 0.3092 / 0.2906 0.1257 / 0.0985 / 0.8612 0.0920 / 0.0784 / 0.8798 0.0898 / 0.0787 / 0.8978
Original VVC VQIR MS-ILLM
[…]
Text+Sketch-25 PerCo-5 DiffEIC-50 RDEIC-2
0.0165 / 0.3409 / 0.0744 0.1256 / 0.0567 / 0.9631 0.0605 / 0.0624 / 0.9664 0.0591 / 0.0485 / 0.9737
Figure 12: More qualitative comparisons between RDEIC and other methods.
Figure 13: More results regarding the balance between smoothness and sharpness.
$\lambda_{s}$=0.0 $\lambda_{s}$=0.1 $\lambda_{s}$=0.2 $\lambda_{s}$=0.3
$\lambda_{s}$=0.4 $\lambda_{s}$=0.5 $\lambda_{s}$=0.6 $\lambda_{s}$=0.7
$\lambda_{s}$=0.8 $\lambda_{s}$=0.9 $\lambda_{s}$=1.0 $\lambda_{s}$=1.1
$\lambda_{s}$=1.2 $\lambda_{s}$=1.3 $\lambda_{s}$=1.4 $\lambda_{s}$=1.5
Figure 14: More results regarding the balance between smoothness and sharpness.
$\lambda_{s}$=0.0 $\lambda_{s}$=0.1 $\lambda_{s}$=0.2 $\lambda_{s}$=0.3
$\lambda_{s}$=0.4 $\lambda_{s}$=0.5 $\lambda_{s}$=0.6 $\lambda_{s}$=0.7
$\lambda_{s}$=0.8 $\lambda_{s}$=0.9 $\lambda_{s}$=1.0 $\lambda_{s}$=1.1
$\lambda_{s}$=1.2 $\lambda_{s}$=1.3 $\lambda_{s}$=1.4 $\lambda_{s}$=1.5
Figure 15: More results regarding the balance between smoothness and sharpness.

### v2

@@ IV-B3 Complexity Comparisons
Figure 7: Ablation studies of the proposed relay residual diffusion and fixed-step fine-tuning on the CLIC2020 dataset.
TABLE II: The impact of RRD and FSFT on performance. BD-Rate is evaluated on CLIC2020 dataset, using DiffEIC [16] as the anchor. DS denotes the number of denoising steps. 2/5 denotes that we use 2 denoising steps for the two models with larger bpp and 5 steps for the remaining models.
Methods DS BD-Rate (%)
DISTS LPIPS PSNR MS-SSIM
DiffEIC 50 0 0 0 0
Baseline 50 1.57 4.88 9.72 3.58
+RRD 2/5 12.63 -4.52 -47.24 -18.65
+RRD+FSFT 2/5 -17.94 -40.79 -61.26 -32.73
V Analysis and Discussions
@@
To provide a more comprehensive analysis of the proposed method, we conduct ablation studies, with the results presented in Fig. 7 and Table II. For the baseline, we employ the same diffusion framework as DiffEIC [16], where the denoising process starts from pure noise. As shown in Fig. 7, our baseline performs similarly to DiffEIC [16], indicating that the choice of compression module has minimal impact on overall performance.
V-A Effectiveness of Relay Residual Diffusion
@@
We first investigate the effectiveness of our proposed relay residual diffusion framework. As shown in Fig. 7 and Table II, by incorporating the proposed relay residual diffusion framework, we achieve better distortion performance and comparable perceptual performance with 2/5 denoising steps compared to the Baseline, which uses 50 denoising steps. The reason behind this is that starting from the compressed latent feature, instead of pure noise, avoids the error accumulation in the initial stage of the denoising process and provides a solid foundation for subsequent detail generation.
Since the time required for the denoising stage is directly proportional to the number of denoising steps, incorporating RRD reduces the denoising time by a factor of 10$\times$ to 25$\times$ compared to the baseline, as shown in Table III.
TABLE III: The impact of RRD on speed. DS denotes the number of denoising steps. FSFT is a fine-tuning strategy that does not affect speed.
Methods DS Denoising Time Speedup
DiffEIC 50 4.342 $\pm$ 0.013 1$\times$
Baseline 50 4.349 $\pm$ 0.013 1$\times$
+RRD 5 0.434 $\pm$ 0.002 10$\times$
2 0.173 $\pm$ 0.001 25$\times$
V-B Analyze of Denoising Steps
@@
Next, we analyze the impact of denoising steps on “Baseline+RRD” to select an appropriate value of $L$ for FSFT strategy. As shown in Fig. 8, for $\lambda_{r}\in\{2,1\}$, the number of denoising steps has minimal effect on compression performance, so that we set $L$ to 2 in this case. For $\lambda_{r}\in\{0.5,0.25,0.1\}$, increasing the denoising steps achieves better perceptual results (lower LPIPS and DISTS values), but leads to degraded fidelity (lower PSNR and MS-SSIM values). To achieve a balance between fidelity and perceptual quality, we set $L$ to 5 here.
Figure 8: The impact of denoising steps on “Baseline+RRD”.
Figure 9: Ablation studies of the diffusion mechanism on CLIC2020 dataset. In the W/o denoising process setting, we train the compression module jointly with the noise estimator but bypass the denoising process during inference. In the W/o diffusion mechanism setting, we train the compression module independently, completely excluding the influence of the diffusion mechanism.
V-C Effectiveness of Fixed-step Fine-tuning
We further demonstrate the effectiveness of the proposed FSFT strategy. As shown in Fig. 7 and Table II, the FSFT strategy significantly improves reconstruction performance across all metrics, indicating that it effectively eliminates the discrepancy between the training and inference phases. Furthermore, as FSFT is a fine-tuning strategy, it does not introduce any additional computational overhead during inference.
@@ V-C Effectiveness of Fixed-step Fine-tuning
Refer to caption Refer to caption
Figure 10: Balancing smoothness versus sharpness. The second and fourth rows show the absolute difference between the reconstructed images and the baseline ($\lambda_{s}=0$).
V-D Smoothness-sharpness Trade-off
To fully leverage the generative potential of pre-trained stable diffusion, we introduce a controllable detail generation method that allows users to explore and customize outputs according to their personal preferences. For this experiment, we used the model trained with $\lambda_{r}=1$. The visualization result is shown in Fig. 10. We control the balance between smoothness and sharpness by adjusting the parameter $\lambda_{s}$, which regulates the amount of high-frequency details introduced into the reconstructed image. Specifically, as the value of $\lambda_{s}$ increases, the image transitions from a smooth appearance to a progressively sharper and more detailed reconstruction.
V-E Role of the Diffusion Mechanism
To further investigate the role of the diffusion mechanism in RDEIC, we design two variants for comparison: 1) W/o denoising process: In this variant, the compression module is trained jointly with the noise estimator, but the denoising process is bypassed during the inference phase. 2) W/o diffusion mechanism: In this variant, the compression module is trained independently, completely excluding the influence of the diffusion mechanism.
As shown in Fig. 9, bypassing the denoising process results in significant degradation, particularly in perceptual quality. This demonstrates that the diffusion mechanism plays a crucial role in enhancing perceptual quality during reconstruction. As shown in Fig. 11, the diffusion mechanism effectively adds realistic and visually pleasing details. By comparing the performance of W/o diffusion mechanism and W/o denoising process in Fig. 9 and Fig. 11, we observe that the compression module trained jointly with the noise estimator outperforms the one trained independently. This demonstrates that the diffusion mechanism also contributes to the compression module.
Fig. 12(a) visualizes an example of bit allocation. It is evident that the model trained jointly with the noise estimator allocates bits more efficiently, assigning fewer bits to flat regions (e.g., the sky in the image). Fig. 12(b) visualizes the cross-correlation between each spatial pixel in $(\boldsymbol{y}-\boldsymbol{\mu})/\boldsymbol{\sigma}$ and its surrounding positions. Specifically, the value at position $(i,j)$ represents cross-correlation between spatial locations $(x,y)$ and $(x+i,y+j)$ along the channel dimension, averaged across all images on Kodak dataset. It is evident that the model trained jointly with the noise estimator exhibits lower latent correlation, suggesting reduced redundancy and more compact feature representations. These results indicate that the diffusion mechanism provides additional guidance for optimizing the compression module during training, enabling it to learn more efficient and compact feature representations.
V-F Limitations
Original W/o diffusion mechanism W/o denoising process RDEIC
Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 11: Impact of diffusion mechanism on reconstruction results.
W/o diffusion W/ diffusion W/o diffusion W/ diffusion
Refer to caption Refer to caption Refer to caption Refer to caption 0.0487 bpp 0.0460 bpp 0.0667 bpp 0.0654 bpp
(a) Bit allocation (b) Cross-correlation
Figure 12: Impact of the diffusion mechanism on the compression module. W/o diffusion denotes the compression module trained independently, while W/ diffusion denotes the compression module trained jointly with the noise estimator. All results are obtained from models trained with $\lambda_{r}=0.5$. (a) An example of bit allocation on the Kodak dataset, with the values normalized for consistency. (b) Latent correlation of $(\boldsymbol{y}-\boldsymbol{\mu})/\boldsymbol{\sigma}$.
Ground Truth 0.0271 bpp Ground Truth 0.0260 bpp
Figure 13: Faces generated at extremely low bitrates.
VI Conclusion